OpenStudy (zeesbrat3):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8.

OpenStudy (jdoe0001):

bearing in mind that the vertex of the parabola is half-way between the directrix and the focus point what do you think are the coordinates of the vertex?

OpenStudy (zeesbrat3):

The vertex would be (2, 0)?

OpenStudy (jdoe0001):

 not quite check.... see what would be the vertex coordinates

OpenStudy (zeesbrat3):

(2,4)

OpenStudy (jdoe0001):

"ith a \(\bf \text{focus at (2, 4)}\) and a directrix of y = 8."

OpenStudy (jdoe0001):

that's just the given focus point but the vertex is half-way between the focus point and the directrix

OpenStudy (zeesbrat3):

Oh! So 8-4 = 4/2 is 2 in each direction. The vertex would be (2,6) with a negative slope since it is opening down?

OpenStudy (jdoe0001):

yeap so now we know the vertex is at (2,6) and the parabola is indeed opening downwards thus noticd the distnace from the vertex to the focus or the directrix, since it is the same 2 units up and 2 units down thus \(\bf (x-{\color{brown}{ h}})^2=4{\color{blue}{ p}}(y-{\color{brown}{ k}}) \\ \quad \\ vertex \ (h,k)\qquad p=\textit{distance from the vertex to focus or directrix} \\ \quad \\ (x-{\color{brown}{ 2}})^2=4({\color{blue}{ 2}})(y-{\color{brown}{ 6}})\)